# Some Fibonacci-Related Sequences

### 22 Dec 2023 v2

@@ 1 Introduction
In addition to studying A105774, we also prove two new results about synchronized sequences, in Theorems 11 and 19.
@@ 5 Guessing and checking
One of the principal tools we use is the “guess-and-check” method. Here we guess a finite automaton computing a sequence based on empirical data, and then use Walnut to verify that it is correct, usually by checking that the defining equation holds. The guessing procedure is described in [9, §5.7].
@@
The numbers $p_{0}(n)$ are precisely those in A004937.
Proof.
We use the following Walnut command:
and Walnut returns TRUE. ∎
Proposition 4.
@@ 6 Inequalities
7 Missing numbers
Let us prove the following result:
Proposition 9.
The natural numbers that do not appear in A105774 are those given by the sequence A004937, that is, $\{\lfloor\varphi^{2}n+{1\over 2}\rfloor\,:\,n\geq 1\}$.
Proof.
We have, using the same idea as in the proof of Proposition 4, that
We can now use Walnut to define a synchronized automaton for A004937:
def a004937 "?msd_fib Ex \$phi2n(2*n,x) & z=(x+1)/2":
The automaton for A004937 is given in Figure 8.
Figure 8: Synchronized Fibonacci DFAO for A004937.
We can now complete the proof as follows:
eval missing "?msd_fib An (n>=1) => ((Ek \$a004937(k,n)) <=> C[n]=@0)":
and Walnut returns TRUE. ∎
@@ 9 Sorting the terms
We just sketch the proof, which depends on linear representations. For more about these, see [9, Chap. 9] and [2]. Given automata for $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$, we can find linear representations counting the number of occurrences of each natural number $n$ in each sequence. Then we can form the linear representation for the difference of these two. Then we use the fact that whether a linear representation is identically 00 is decidable. ∎
@@ 13 Other results
For the remainder of this section, let $a(n)$ represent (as before) the $n$’th term of A105774, let $b(n)=\lfloor n\varphi\rfloor$, and let $c(n)=\lfloor n\varphi^{2}+1/2\rfloor$. We will find some relations concerning the composition of the functions $a$, $b$, and $c$.
@@
    For $n\geq 1$, the difference $a(b(n))-b(a(n))$ is exactly the parity of the number of $1$’s in the Fibonacci representation of $n-1$, that is, sequence A095076.
@@
eval test1 "?msd_fib An \$abmba(n+1,0) <=> \$even1(n)":
@@ 14 Going further
It is possible to study many similar recurrences using exactly the same techniques. We list just two possibilities; many more should occur to the interested reader.

### 01 Jan 2024 v3

@@ 1 Introduction
In addition to studying A105774, we also prove two new results about synchronized sequences, in Theorems 10 and 18.
@@ 5 Guessing and checking
One of the principal tools we use is the “guess-and-check” method. Here we guess a finite automaton computing a sequence based on empirical data, and then use Walnut to verify that it is correct, usually by checking that the defining equation holds. The guessing procedure is described in [10, §5.7].
@@
and Walnut returns TRUE. ∎
Finally, we prove the result for $p_{0}$:
Proposition 5.
The natural numbers $p_{0}(n)$, i.e., the numbers that do not appear in A105774 are those given by the sequence A004937, that is, $\{\lfloor\varphi^{2}n+{1\over 2}\rfloor\,:\,n\geq 1\}$.
Proof.
We have, using the same idea as in the proof of Proposition 3, that
We can now use Walnut to define a synchronized automaton for A004937:
def a004937 "?msd_fib Ex \$phi2n(2*n,x) & z=(x+1)/2":
The automaton for A004937 is given in Figure 7.
Figure 7: Synchronized Fibonacci DFAO for A004937.
We can now complete the proof as follows:
@@ 8 Sorting the terms
We just sketch the proof, which depends on linear representations. For more about these, see [10, Chap. 9] and [2]. Given automata for $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$, we can find linear representations counting the number of occurrences of each natural number $n$ in each sequence. Then we can form the linear representation for the difference of these two. Then we use the fact that whether a linear representation is identically 00 is decidable. ∎
@@ 12 Other results
For the remainder of this section, let $a(n)$ represent (as before) the $n$’th term of A105774, let $b(n)=\lfloor n\varphi\rfloor$, and let $c(n)=\lfloor n\varphi^{2}+1/2\rfloor$. We will find some relations concerning the composition of the functions $a$, $b$, and $c$. We introduce the function $x(n):=a(b(n))-b(a(n))$.
@@
    For $n\geq 1$, the difference $x(n):=a(b(n))-b(a(n))$ is exactly the parity of the number of $1$’s in the Fibonacci representation of $n-1$, that is, sequence A095076.
@@
eval test1 "?msd_fib An \$xx(n+1,0) <=> \$even1(n)":
@@
13 Carlitz-style results
[…]
We adopt the following notation: if $u=u_{1}\cdots u_{t}\in\{a,b,d,x\}^{*}$, then by $u(n)$ we mean $(u_{1}\circ\cdots\circ u_{t})(n)$. Recall that $x$ was defined in Section 12.
Theorem 23.
Let $u\in\{b,d\}^{+}$ be a nonempty word with $i$ copies of $b$ and $j$ copies of $d$. Then for all $n\geq 1$ we have
$au(n)=F_{i+2j}\cdot ab(n)+F_{i+2j-1}\cdot a(n)+C(u)\cdot(2x(n)-1),$
where $C(u)$ is defined as follows:
$C(u)=\begin{cases}0,&\text{if $u=b$};\\ 1,&\text{if $u=d$};\\ F_{i+2j-1}+C(v),&\text{if $u=vb$};\\ F_{i+2j-1}-C(v),&\text{if $u=vd$}.\end{cases}$ (4)
Proof.
First we need some lemmas:
Lemma 24.
We have
  • (a)
    $xb(n)=x(n)$ for all $n\geq 0$.
  • (b)
    $xd(n)=1-x(n)$ for all $n\geq 1$.
  • (c)
    $abb(n)=ab(n)+a(n)+2x(n)-1$ for all $n\geq 1$.
  • (d)
    $abb(n)=ad(n)$ for all $n\geq 0$.
  • (e)
    $abd(n)=2ab(n)+a(n)+2x(n)-1$ for all $n\geq 1$.
Proof.
These can be proved with Walnut as follows. Recall that phin is Walnut’s name for $b$ and phi2n is Walnut’s name for $d$.
eval checka "?msd_fib An,y,z,w (n>=0 & \$xx(n,y) & \$phin(n,z) & \$xx(z,w))
   => w=y":
eval checkb "?msd_fib An,y,z,w (n>=1 & \$xx(n,y) & \$phi2n(n,z) & \$xx(z,w))
   => w+y=1":
eval checkc "?msd_fib An,y,z,w,t (n>=1 & \$abb(n,y) & \$ab(n,z) &
   \$a105774(n,w) & \$xx(n,t)) => y+1=z+w+2*t":
def ad "?msd_fib Ew \$phi2n(n,w) & \$a105774(w,z)":
def abd "?msd_fib Ew,y \$phi2n(n,w) & \$phin(w,y) & \$a105774(y,z)":
eval checkd "?msd_fib An,z (n>=0 & \$ad(n,z)) => \$abb(n,z)":
eval checke "?msd_fib An,y,z,w,t (n>=1 & \$abd(n,y) & \$ab(n,z) &
   \$a105774(n,w) & \$xx(n,t)) => y+1=2*z+w+2*t":
and Walnut returns TRUE for all. ∎
Now we can prove the desired result by induction on $|u|$. It is clear from the above results for $|u|=1$. Now assume $|u|\geq 2$ and consider two cases: $u=vb$ and $u=vd$.
$u=vb$: In this case we have, by induction, that
$av(n)=F_{i+2j-1}\cdot ab(n)+F_{i+2j-2}\cdot a(n)+C(v)\cdot(2x(n)-1).$
Substituting $b(n)$ for $n$, we get
$avb(n)$ $=F_{i+2j-1}\cdot abb(n)+F_{i+2j-2}\cdot ab(n)+C(v)\cdot(2xb(n)-1)$
$=F_{i+2j-1}\cdot(ab(n)+a(n)+2x(n)-1)+F_{i+2j-2}ab(n)+C(v)\cdot(2xb(n)-1)$
$=F_{i+2j}\cdot ab(n)+F_{i+2j-1}a(n)+(C(v)+F_{i+2j-1})(2x(n)-1),$
as desired.
$u=vd$: By induction we have
$av(n)=F_{i+2j-2}\cdot ab(n)+F_{i+2j-3}\cdot a(n)+C(v)\cdot(2x(n)-1).$
Substituting $d(n)$ for $n$, we get
$avd(n)$ $=F_{i+2j-2}\cdot abd(n)+F_{i+2j-3}\cdot ad(n)+C(v)\cdot(2xd(n)-1)$
$=F_{i+2j-2}\cdot(2ab(n)+a(n)+2x(n)-1)+F_{i+2j-3}(ab(n)+a(n)+2x(n)-1)$
$\quad+C(v)(2(1-x(n))-1)$
$=F_{i+2j}\cdot ab(n)+F_{i+2j-1}a(n)+(F_{i+2j-1}-C(v))(2x(n)-1).$
This completes the proof. ∎
We can obtain a linear representation for $C(u)$, as follows.
Corollary 25.
Define vectors $L,R$ and a matrix-valued morphism $\mu$ as follows:
$L$ $=\left[\begin{array}[]{ccc}0&0&1\end{array}\right]$
$\mu(b)$ $=\left[\begin{array}[]{ccc}0&1&0\\ 1&1&0\\ 0&1&1\end{array}\right]$
$\mu(d)$ $=\left[\begin{array}[]{ccc}0&0&1\\ 0&1&1\\ 1&2&1\end{array}\right]$
$R$ $=\left[\begin{array}[]{c}1\\ 0\\ 0\\ \end{array}\right].$
Then $C(u)=L\mu(u)R$.
Proof.
We first obtain the following relations, for $v\in\{b,d\}^{+}$:
$C(vbb)$ $=C(v)+C(vb)+C(vd)$
$C(vbd)$ $=C(vd)$
$C(vdb)$ $=C(vb)+2C(vd)$
$C(vdd)$ $=C(v)+C(vb)+C(vd).$
We just show how to obtain the first relation; the others are similar. Letting $u=vbb$, and using (4), we get
$C(u)$ $=C(vbb)=F_{i+2j-1}+C(vb)$
$C(vb)$ $=F_{i+2j-2}+C(v)$
from which it follows that $C(u)=F_{i+2j}+C(v)$. On the other hand, from (4), we have $C(vd)=F_{i+2j-1}-C(v)$ and so
$C(v)+C(vb)+C(vd)=C(v)+F_{i+2j-2}+C(v)+F_{i+2j-1}-C(v)=C(v)+F_{i+2j},$
and the first identity follows.
Once we have the relations, we can combine them in the usual way to obtain the given linear representation. ∎
@@ 14 Going further
It is possible to study many similar recurrences using exactly the same techniques. We list just three possibilities; many more should occur to the interested reader.
@@
    One can study a variation of the defining recurrence (1) where the Lucas numbers replace the Fibonacci numbers. The resulting sequence is A118287 in the OEIS, and is computed by a $102$-state synchronized automaton.
    With this automaton one can easily prove observations such as
    A118287$(n)$ = A118287$(n-2)$ if and only if $n$ belongs to A003231.
  2. 2.